# Gravitational Waves from Massive Black Hole Mergers in ASTRID: Predictions for LISA

### v1

@@ 4.2 MBH population detected by LISA
In this section, we investigate the MBH population involved in the mergers detected by LISA. In Fig. 12, we show the PDF for the ratio of the merging MBH mass $M_{\mathrm{tot}}=M_{\mathrm{1}}+M_{2}$ to the total MBH mass in the host galaxy (left), the Eddington ratio $f_{\mathrm{Edd}}$ (middle), and the bolometric luminosity $L_{\mathrm{bol}}$ (right). In each panel, we plot the detectable sources with SNR larger than 10 (histograms) and with SNR $>1000$ (curves) for $z=1.5$ (blue) and $z=0.8$, respectively.
@@
From the distribution of $M_{\mathrm{tot}}/M_{\mathrm{BH,gal}}$, we can see that in roughly half of host galaxies, merging binaries constitute over 90% of the total black hole mass. This means that the remnant MBHs are typically the galaxy central black hole. The distribution is similar at $z=0.8$ and $z=1.5$, except that at $z=1.5$, there are more sources with masses over $90\%$ of the total BH mass, while there are more remnant MBHs which have mass between $0.8\sim 0.9\,M_{\mathrm{BH,gal}}$ at $z=0.8$. This can be explained by the larger galaxies at lower redshift, which host many more small-mass wandering MBHs.
We then investigate the accretion efficiency of the merging MBHs. The peak of the Eddington ratio $f_{\mathrm{Edd}}=L_{\mathrm{bol}}/L_{\mathrm{edd}}$ the $f_{\mathrm{edd}}$ distribution is skewed to lower values at lower redshift: at $z=1.5$ the peak is $\log(f_{\mathrm{edd}})=-1.75$ which evolves to $\log(f_{\mathrm{edd}})=-2$ by $z=0.8$.
@@
Unlike the Eddington ratio, $L_{\mathrm{bol}}$ peaks around $\log(L_{\mathrm{bol}})\sim 42.5\,\mathrm{erg/s}$ for both $z=0.8$ and $z=1.5$, while the high-$z$ distribution has a larger low-luminosity tail. Defining the MBHs with $L_{\mathrm{bol}}>10^{43}\ \mathrm{erg/s}$ as AGN (Chen et al., 2023) (marked by the vertical black dotted line in the right panel in Fig. 12) we find that only $10.6\%$ of detectable (i.e., with SNR larger than 10) sources involve AGN activity at $z=1.5$, and $16.6\%$ at $z=0.8$. Thus only a small fraction of LISA sources will have detectable electromagnetic counterparts. We check that most of these mergers involving the AGNs are from massive mergers with $M_{\mathrm{tot}}\gtrsim 10^{6.5}$ $\mathrm{M}_{\odot}$. Recently, Dong-Páez et al. (2023) also investigated the AGN signatures of LISA sources, but at a higher redshift of $z\sim 3.5$, and found $4\sim 20\%$ of sources to be detectable through AGN signatures. Our results are in general agreements despite the different source redshift distributions.
5 Discussion
5.1 The Influence of Eccentricity
As we show in Fig. 7 and 8, including the effect of eccentricity (the Ecc run) enhances the detectabilit of events with LISA. The detectable mass range expands from $10^{8}$ $\mathrm{M}_{\odot}$ to $10^{9}$ $\mathrm{M}_{\odot}$, and the detection rate at low redshift ($z<0.5$) and at high mass ($M_{\mathrm{BH}}>10^{7.5}$ $\mathrm{M}_{\odot}$) increases by an order of magnitude. In this section we discuss in detail how the eccentricity makes such a difference.
In the period a few years before coalescence, the eccentricity of a binary will rapidly reduce via GW emission driven circularization. If this circularization process causes the eccentricity to reach close to zero, or the dominant harmonic mode becomes close to $n=2$, we find that the behavior of the Ecc run is very similar to that of the Circ run (shown in the left and middle panels of Fig. 2). Since the total SNR is typically dominated by the emission in the short phase before coalescence, these fully circularized sources have similar detectability in both the Ecc and Circ runs. The difference in the detectable population for Ecc and Circ runs thus mainly comes from mergers with large residual eccentricity.
One caveat about the Ecc run is that we directly adopt the orbital eccentricity measured from ASTRID as the value at $t_{\mathrm{start}}$, i.e., the initial eccentricity observed by LISA. The hardening timescale for binaries to evolve from $\sim 1$ kpc into the GW-emitting regime could be over $1$ Gyr (Chen et al., 2020; Kelley et al., 2017), and the eccentricity evolution during this process is dependent on the local galaxy environment. As shown by Sesana (2010), stellar dynamics causes a dramatic increase in orbital eccentricity, especially for initially already mildly eccentric and/or unequal mass binaries. In the MAGICS simulation suite (Chen et al., 2024; Zhou et al., 2025b; Mukherjee et al., 2025), the authors found that the orbital eccentricities increase during the evolution of the bound binary, but are constant once hard binaries form. Romero-Shaw et al. (2024) showed that binaries driven by the circumbinary disc will evolve to an equilibrium eccentricity: $0.2\lesssim e_{\mathrm{eq}}\lesssim 0.5$. More importantly, when the binary is observed at $t_{\mathrm{start}}$, it is likely that it has gone through GW circularization since the GW emission timescale is typically over 100 years. For the observed merger sources for which $t_{\mathrm{obs}}$ is within 4 years of the merger and which have rapidly decreasing eccentricity, circularization becomes the dominant effect. Considering that most of our detected sources are merger sources (see Fig. 8), the LISA sources in our current work are therefore those with higher eccentricities. In future work, we plan to implement an eccentricity evolution model, which depends on the local environment of each binary, such as the stellar profile index and density, to track the eccentricity evolution of the unbound binaries in a more self-consistent way.
5.2 Binary Delayed Timescale
In this work, we do not apply any binary delay model to account for the evolution timescale from a $1$ kpc separation (i.e., the resolution limit of ASTRID) to the GW regime. Considering unresolved dynamical friction and stellar scattering, Chen et al. (2022b) estimated the hardening timescale for the MBH binary population in ASTRID with $z>3$. Incorporating the delay time, it was found that the decrease in the resultant detection rate is smaller than one order of magnitude. Katz et al. (2020) evolved the MBH merger population in Illustris by post-processing using the treatments of Kelley et al. (2017) and Dosopoulou & Antonini (2017), and found that the detection rate only drops by a maximum of half. Salcido et al. (2016) came to a similar conclusion: the merger rate is similar across all hardening models, and their impact is much smaller than the impact of changing the black hole seed mass adopted. These previous analyses imply that including the binary delay may only change in a minor fashion for the predicted detection rate plotted in Fig. 8.
@@ 6 Conclusion
In this work, we estimate the GW signal emitted by MBH mergers down to $z=0$ in the cosmological simulation ASTRID, and make predictions for the LISA detection rate.
@@
Using the recorded trajectories of MBHs in ASTRID, we are able to measure the orbital eccentricity $e$ of merging objects. We find that the MBH mergers in general have high eccentricity: $e$ peaks at $\gtrsim 0.8$ across the whole redshift range. It evolves slightly to smaller values at low redshift: from $0.88$ at $z>3$ to $0.80$ at $z<0.5$, indicating that a larger fraction of binaries at low redshift merge with circular orbits. This can be explained by the shorter hardening timescales for binaries with high eccentricity (Chen et al., 2022b). To give a better description of the $e$ distribution, we provide a fitted $e$ PDF function (Equation 9), which includes the redshift evolution.
@@
To investigate the influence of eccentricity on GW emission in the LISA band ($110^{-5}<f<$ 1 Hz) we generate two sets of signals: a Circ run that assumes circular orbits for all the sources, and a Ecc run that includes the effect of eccentricity in the calculation of GW waveforms. Given the black hole seed masses in ASTRID of $M_{\mathrm{BH,seed}}\sim 5\times 10^{4}$ $\mathrm{M}_{\odot}$, we find that LISA will probe MBH mergers across a wide parameter space, covering objects with $M_{\mathrm{BH}}\lesssim 10^{9}$ $\mathrm{M}_{\odot}$ and $z\lesssim 8$.
@@
We implement a Monte Carlo analysis technique to estimate the LISA detection rate. This enables us to generate multiple realizations of the merger population detectable by LISA, and avoids the assumption of a specific observation time for each MBH binary. We can therefore consider the signals emitted from different binary evolution phases, i.e., from both inspiral sources and merging sources. We generate 10,000 realizations of the merging MBH populations that would be detected by a 4-year LISA observation. Our prediction is averaged over these realizations: The total LISA detection rate predicted by ASTRID is $5.6$/yr for the Circ run and $10.5$/yr for the Ecc run. Among them, 46% (4.8 events per year) are inspiral sources for Ecc, and only $0.3\%$ ($\sim 10^{-4}$ event per year) are for Circ.
@@
While the detectable parameter space for the eccentric and circular binaries largely overlaps, including the eccentricity will expand the detectable mass range by one order of magnitude: from $M_{\mathrm{BH}}\sim 10^{8}$ $\mathrm{M}_{\odot}$ to $\sim 10^{9}$ $\mathrm{M}_{\odot}$. The redshift peak for the LISA sources also shifts to a lower value in the Ecc run ($z_{\rm peak}=0.8$ compared to the Circ run ($z_{\rm peak}$ of $z=1.9$). At low redshift ($z<0.5$) and high BH mass ($M_{\mathrm{bh}}>10^{7.5}$ $\mathrm{M}_{\odot}$), the Ecc run predicts a detection rate one order of magnitude higher than Circ run. The improvement in detectability comes from the inspiral sources, which coalesce after the 4-year LISA observation phase. The residual eccentricity in the inspiral phase produces waveforms covering a wide frequency range and GW signals better matched to the LISA high-sensitivity band. Our results underscore the importance of incorporating the orbital eccentricity in the predictions for upcoming GW detection, which requires accurately modeling the dynamics of MBH binaries and their merger histories.
@@
We also investigate the host galaxies of detectable mergers. Unlike the PTA CW sources, the LISA detectable mergers will be observed to lie in various galaxy environments, covering a wide range of mass ($10^{7}-10^{11}$ $\mathrm{M}_{\odot}$), and SFR ($10^{-3}-10^{2}$ $\mathrm{M}_{\odot}$/yr). Notably, a significant fraction ($\sim 20-30\%$) of the detectable sources occur in satellite galaxies, sometimes even outside the virial radius. The merger hosts are also much more highly star forming than the general population, with a mean SFR fifty times higher for a given galaxy mass. The relationship between SFR and galaxy mass for merging MBHs is tight, tracing a distinct population, and this sequence is well described by a power law where SFR= $10^{-8}M_{\rm gal}^{0.9}$ $\mathrm{M}_{\odot}$/yr, with a scatter of 0.17 dex. Most of the remnant MBHs are the central MBH of the galaxy, while only a small fraction (about $10\%$) are AGNs with a bolometric luminosity larger than $10^{43}$ erg/s. This implies that it will be difficult to search for the electromagnetic counterpart of LISA sources.
@@ Acknowledgements
BYW thanks Yili Han and Handan Zhang for their helpful advice on the visualization. TDM acknowledges funding from NASA ATP 80NSSC20K0519, NSF PHY-2020295, NASA ATP NNX17AK56G, and NASA ATP 80NSSC18K101, NASA Theory grant 80NSSC22K072. NC acknowledges support from the Schmidt Futures Fund. YN acknowledges support from the ITC Postdoctoral Fellowship. SB acknowledges funding from NASA ATP 80NSSC22K1897. ASTRID was run on the Frontera facility at the Texas Advanced Computing Center.

### v2

@@ 4.2 MBH population detected by LISA
In this section, we investigate the MBH population involved in the mergers detected by LISA. In Fig. 12, we show the PDF for the ratio of the merging MBH mass $M_{\mathrm{tot}}=M_{\mathrm{1}}+M_{2}$ to the total MBH mass in the host galaxy (left), the bolometric luminosity $L_{\mathrm{bol}}$ (middle), and the Eddington ratio $f_{\mathrm{edd}}$ (right). In each frame, besides showing all the detectable mergers in the Ecc across the whole redshift range using the grey histogram, we also plot the distribution of MBHs involving mergers with $M_{\mathrm{tot}}>10^{6}$ $\mathrm{M}_{\odot}$ (black solid curve) and $M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$ (red solid curve). The dashed histograms correspond to the mergers occurring at high redshift ($z\geq 2$) with $M_{\mathrm{tot}}>10^{6}$ $\mathrm{M}_{\odot}$ (black) or $M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$ (red). $M_{\mathrm{tot}}>10^{6}$ $\mathrm{M}_{\odot}$ is where the detection rate peaks in the right panel of Fig. 8, so the black solid curve represents most of the mergers, and the gap between the black solid curve and the grey histogram is the contribution from the low-mass mergers.
@@
From the distribution of $M_{\mathrm{tot}}/M_{\mathrm{BH,gal}}$, we can see that in most host galaxies (85%) merging binaries constitute over half of the total black hole mass, which means the remnant MBHs are typically the galaxy’s central black hole. The distribution is similar among the whole population and the high-redshift mergers, while for mergers at the high-mass end ($M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$), a larger fraction of the remnant MBHs (96%) dominates the galaxy black hole mass with $M_{\mathrm{tot}}/M_{\mathrm{BH,gal}}>0.5$.
@@
We then investigate the accretion of the merging MBHs. For all the detectable mergers in Ecc and the population with $M_{\mathrm{tot}}>10^{6}$ $\mathrm{M}_{\odot}$, the $L_{\mathrm{bol}}$ distribution both peak at $\log(L_{\mathrm{bol}})=42.6\,\mathrm{erg/s}$. 70% of total mergers have $\log(L_{\mathrm{bol}})>42$ erg/s. As expected, massive mergers in general have higher luminosity: $L_{\mathrm{bol}}$ peaks at $10^{44.3}$ erg/s for mergers with $M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$. Only a small fraction of mergers have relatively high luminosity: less than $1\%$ detected mergers have luminosity larger than $10^{44}$ erg/s, and 66% of them are from merger with $M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$. Recently, Dong-Páez et al. (2023) also investigated the AGN signatures of LISA sources, but at a higher redshift of $z\sim 3.5$, and found $4\sim 20\%$ of sources to be detectable through AGN signatures. Our results are in general agreements despite the different source redshift distributions.
For the Eddington ratio $f_{\rm edd}$, there is only a minor distinction between the population with different mass cuts, while it presents evolution across the redshift: for both the merger with $M_{\rm tot}\geq 10^{6}$ $\mathrm{M}_{\odot}$ and $M_{\rm tot}\geq 10^{8}$ $\mathrm{M}_{\odot}$ over the whole redshift range, $f_{\rm edd}$ peaks at $0.01$; while for $z\geq 2$, they have higher Eddington ratio, both peaking at $0.02$. Note that $f_{\rm edd}$ can be larger than 1 since we adopt a super-Eddington accretion.
5 Discussion: Binary Delay Timescale
In this work, we do not apply any binary delay model to account for the evolution timescale from a separation of $1$ kpc to the GW regime.
This choice is motivated by the fact that ASTRID already includes the effects of DF, which in previous works have been shown to dominate delays in MBH binary dynamics (prior to the GW-driven phase). In particular, Li et al. (2022) evolved unbound MBH pairs in the TNG50-3 cosmological simulation from kpc scales to coalescence using a detailed postprocessing formalism that incorporates DF, loss-cone scattering, viscous drag from a circumbinary disk, and GW emission. Their results showed that the DF-dominated phase accounts for the largest fraction of the total binary hardening timescale.
Several previous studies incorporated binary delay based on postprocessing the simulation binaries with analytical prescriptions to predict MBH merger rates and associated LISA event rates. Katz et al. (2020) post-processed the MBH merger population from the Illustris cosmological simulation using the prescriptions of Kelley et al. (2017) and Dosopoulou & Antonini (2017). They found that considering the delay timescale only reduced the detection rate by at most a factor of two. Similarly, Salcido et al. (2016) concluded that merger rates are relatively insensitive to the specific adopted hardening model, and the choice of MBH seed mass has a much stronger influence. Volonteri et al. (2020) directly compared the binary population subject only to DF with those incorporating additional hardening mechanisms, including stellar scattering and gas torques. Those authors also found that these additional effects have a minor impact on the predicted detection rate, significantly smaller than that of DF delay. Together, these findings suggest that incorporating a binary delay model would only modestly alter the predicted detection rate shown in Fig. 8.
Self-consistently tracking the binary dynamics on small scales requires simulations with higher resolution. The MAGICS simulation suite (Chen et al., 2024; Zhou et al., 2025b; Mukherjee et al., 2025) resimulates 15 merging systems extracted from ASTRID with the resolution up to 62.5 $\mathrm{M}_{\odot}$, and followed the evolution from $10$ kpc separation into hard binary phase ($\sim 10^{-3}$ pc). These works also give a more complete description of the orbital eccentricity evolution: they found after the binaries are gravitationally bound, their eccentricity remains nearly the same for those with $e_{0}<0.6$; while goes through a slight increase for mergers with $e_{0}>0.6$. Although by now MAGICS has only investigated a small sample out of the merger population in ASTRID, upcoming works will include merging systems covering a larger parameter space. Simulations like MAGICS that directly resolve the binary formation and evolution will continue to provide valuable insights to evaluate and calibrate the results from the large-volume simulations.
@@ 6 Conclusion
In this work, we estimate the GW signal emitted by MBH mergers down to $z=0$ in the cosmological simulation ASTRID, and make predictions for the LISA detection rate. Using the recorded trajectories of MBHs in ASTRID, we are able to measure the orbital eccentricity $e_{0}$ of the binaries at scales of $\sim 1\ \rm{kpc}$. We find that the MBH mergers in general have high eccentricity: $e_{0}$ peaks at $\gtrsim 0.8$ across the whole redshift range. It evolves slightly to smaller values at low redshifts: from $0.88$ at $z>3$ to $0.80$ at $z<0.5$, indicating that a larger fraction of binaries at low redshifts merge with circular orbits. This can be explained by the shorter hardening timescales for binaries with high eccentricity (Chen et al., 2022b). To give a better description of the $e_{0}$ distribution, we provide a fitted $e_{0}$ PDF function (Equation 9), which includes the redshift evolution.
@@
To investigate the influence of eccentricity on GW emission in the LISA band ($10^{-5}\,{\rm Hz}<f<1\,\rm Hz$) we generate two sets of signals: a Circ that assumes circular orbits for all the sources, and a Ecc that includes the effect of eccentricity in the calculation of GW waveforms. Given the black hole seed masses in ASTRID of $M_{\mathrm{BH,seed}}\sim 5\times 10^{4}$ $\mathrm{M}_{\odot}$ and a SNR threshold of 10, we find that LISA will probe MBH mergers across a wide parameter space, covering mergers with $M_{\mathrm{BH}}\lesssim 10^{9}$ $\mathrm{M}_{\odot}$ and $z\lesssim 8$.
@@
We implement a Monte Carlo analysis technique to estimate the LISA detection rate. This enables us to generate multiple realizations of the merger population detectable by LISA within the 4-year observation, and avoids the assumption of a specific observation time for each MBH binary. We can therefore consider the signals emitted from different binary evolution phases, i.e., from both inspiral sources and merging sources. We generate 10,000 realizations of the merging MBH populations that would be detected by a 4-year LISA observation. Our prediction is averaged over these realizations: the total LISA detection rate predicted by ASTRID is $5.6$/yr for the Circ and $10.5$/yr for the Ecc. Among them, 46% (4.8 events per year) are inspiral sources for Ecc, and only $0.3\%$ ($\sim 10^{-4}$ event per year) are for Circ. Our predicted number of detections is noticeably higher than those found in previous studies based on large-volume hydrodynamical simulations such as EAGLE (Salcido et al., 2016) and Illustris (Katz et al., 2020), which reported rates of 0.5similar-toabsent0.5$\sim$ 0.5∼ 0.52222 per year. This also places our results closer to those from semi-analytical models, which have generally predicted higher detection rates ( 8similar-toabsent8$\sim$ 8∼ 825252525 per year) (Bonetti et al., 2019; Berti et al., 2016; Klein et al., 2016; Sesana et al., 2011; Arun et al., 2009; Izquierdo-Villalba et al., 2024), suggesting a convergence between these two modeling approaches.
@@
Including the eccentricity will expand the detected mass range by one order of magnitude: from $M_{\mathrm{BH}}\sim 10^{8}$ $\mathrm{M}_{\odot}$ in the circular scenario to $\sim 10^{9}$ $\mathrm{M}_{\odot}$, and shift the peaks of redshift of LISA sources to a lower value: $z_{\rm peak}=0.8$ in Ecc compared to the $z_{\rm peak}=1.9$ in Circ. At low redshift ($z<0.5$) and high mass end ($M_{\mathrm{tot}}>10^{7.5}$ $\mathrm{M}_{\odot}$), the Ecc predicts a detection rate one order of magnitude higher than Circ. The improvement in detectability comes mainly from the inspiral sources, which coalesce after the 4-year LISA observation phase. The residual eccentricity in the inspiral phase produces waveforms covering a wide frequency range and GW signals better matching the LISA high-sensitivity band. Our results underscore the importance of incorporating the orbital eccentricity in the predictions for upcoming GW detection, which requires accurately modeling the dynamics of MBH binaries.
@@
We also investigate the host galaxies of detectable mergers. Unlike the PTA CW sources, LISA detectable mergers will be observed to lie in various galaxy environments, covering a wide range of mass ($10^{7}-10^{11}$ $\mathrm{M}_{\odot}$), and SFR ($10^{-3}-10^{3}$ $\mathrm{M}_{\odot}$/yr). Notably, a significant fraction ($\sim 20-30\%$) of the detectable sources occur in satellite galaxies, sometimes even outside the virial radius. Most of the remnant MBHs are the most massive MBH of the host galaxy. For all the detectable mergers in Ecc, the distribution of bolometric luminosity peaks at $L_{\rm bol}=10^{42.6}\ \rm erg/s$. Massive merger with $M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$ in general have larger luminosity, peaking at $L_{\rm bol}=10^{44.3}\ \rm erg/s$. Only 1% of detectable mergers have luminosity larger than $10^{44}\ \rm erg/s$, 66% of which are mergers with $M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$.
@@ Acknowledgements
YZ acknowledges helpful discussion with Lei Hu. BYW thanks Yili Han and Handan Zhang for their helpful advice on the visualization. TDM acknowledges funding from NASA ATP 80NSSC20K0519, NSF PHY-2020295, NASA ATP NNX17AK56G, and NASA ATP 80NSSC18K101, NASA Theory grant 80NSSC22K072. NC acknowledges support from the Schmidt Futures Fund. YN acknowledges support from the ITC Postdoctoral Fellowship. SB acknowledges funding from NASA ATP 80NSSC22K1897. ASTRID was run on the Frontera facility at the Texas Advanced Computing Center.